# Multi-Agent Large Language Models for Zero-Shot Robotics Manipulation

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable planning abilities across various domains, including robotics manipulation and navigation. While recent efforts in robotics have leveraged LLMs both for high-level and low-level planning, these approaches often face significant challenges, such as hallucinations in long-horizon tasks and limited adaptability due to the generation of plans in a single pass without real-time feedback. To address these limitations, we propose a novel multi-agent LLM framework, Multi-Agent Large Language Model for Manipulation (MALMM) that distributes high-level planning and low-level control code generation across specialized LLM agents, supervised by an additional agent that dynamically manages transitions. By incorporating observations from the environment after each step, our framework effectively handles intermediate failures and enables adaptive re-planning. Unlike existing methods, our approach does not rely on pre-trained skill policies or in-context learning examples and generalizes to a variety of new tasks. We evaluate our approach on nine RLBench tasks, including long-horizon tasks, and demonstrate its ability to solve robotics manipulation in a zero-shot setting, thereby overcoming key limitations of existing LLM-based manipulation methods.
@@ I INTRODUCTION
Robotic manipulation has seen impressive advancements, enabling agents to handle increasingly complex tasks with greater precision and efficiency [1, 2]. Current solutions, however, often struggle with generalization, particularly when trained using imitation learning [3, 4]. Agents trained in this manner typically excel at specific tasks but lack the adaptability to handle new tasks. One major drawback of imitation learning is the labor-intensive and time-consuming process for data collection [5], which limits the scalability of these systems. Moreover, training agents using policy-based methods for even a single task requires thousands of episodes [6], making the approach computationally expensive and inefficient. Effective manipulation agents, however, should demonstrate a deeper understanding of their environment [7]. This involves not only recognizing and grounding relevant objects but also understanding the relationships between them [8]. Equipped with this knowledge, agents can then plan and execute actions more efficiently, adapting to changes in their environment or task requirements.
@@
Recent advancements in Large Language Models (LLMs) have demonstrated remarkable generalization and reasoning capabilities across diverse domains such as commonsense [9], mathematical [10], and symbolic reasoning [11]. These models, particularly when scaled to billions of parameters [12], exhibit emergent abilities to break down complex tasks into simpler steps through structured reasoning techniques like “chain of thought” [13] and prompting [14]. LLMs have already shown promise in high-level planning tasks across fields, suggesting their potential for solving similar challenges in robotics manipulation by offering more flexible, generalizable solutions. This opens new possibilities for improving manipulation systems beyond their current limitations.
Recent LLM-based approaches to robotics planning have shown promise but also face challenges. A major issue is the tendency of LLMs [15] to produce hallucinations, especially in long-horizon tasks where precision and adaptability are essential. This issue arises from their difficulty in managing lengthy conversations which is common in long-horizon manipulation tasks [16]. As a result, LLMs may disregard geometric constraints, parameters of predefined functions, and can even forget the goal. Additionally, most LLM-based robotics manipulation methods [17, 18] generate entire plans in a single pass without incorporating environmental feedback during execution. This limitation prevents methods to dynamically adjust or replann in response to unexpected outcomes or intermediate failures.
@@
To mitigate these issues, we here propose a Multi-Agent Large Language Model for Manipulation (MALMM) that leverage the collective intelligence and the specialized skills of multiple agents for complex manipulation tasks. Our approach incorporates agents dedicated to high-level planning, low-level code generation and a supervisor that oversees transitions between other agents and tools. We show that through the use of multiple specialized agents in a multi-agent setup, we are able to mitigate the hallucination issues observed in the case of a single agent, see Fig. 1. Our framework further integrates environmental observations at each execution step and can adapt to unforeseen changes of the environment.
In summary, we propose the following contributions:
We introduce the first multi-agent LLM framework for robotics manipulation that can generalize to unseen tasks without training or in-context learning examples.
Our framework integrates intermediate environmental observations and enables dynamic re-planning in response to unforeseen changes of the environment.
Our experiments demonstrate significant improvements of MALMM over the state of the art and show the benefit of a multi-agent framework through extensive ablation studies.
@@ II-A Language Grounded Robotics
Language instructions for robotics serve as a tool for compositional goal specification [19] and allow compositionally scalable generalization to new tasks [20]. The literature around language grounding is vast, ranging from classical tools such as lexical analysis, formal logic, and graphical models to interpreting language instructions [19, 21, 22, 23]. Recently, there has been a lot of effort to utilize the impressive compositional understanding capability of Large Language Models to perform language grounding for robotics [20, 24, 25, 26]. Additionally, recent advancements use pre-trained large language models (LLMs) for their open-world knowledge, tackling more challenging tasks such as 3D robotic manipulation and leveraging code generation capabilities to produce high-level, semantically rich procedures for robot control.
@@ II-B LLM for Robotics
Language Models have been used for diverse applications in the field of robotics ranging from a reward function [27, 28] and data augmentation [29] for policy training, for task planning [17, 30, 31, 32], failure summarization to guide language based planner [33], and policy program synthesis [18]. VoxPoser [34] and Language to Rewards [35] uses LLM for generating reward regions for assisting external trajectory optimizers to compute trajectory. Our work relates to studies where LLM is used for generating plans for completing manipulation tasks. Most of these prior works [36, 32, 30, 34] relied on pre-trained skills, motion primitives and trajectory optimizers and focused primarily on high-level planning. The approach in [17] is closest to our work, it deploys a language model to generate high-level plans and uses its code generation capabilities to convert those high-level plans into low-level control. However, LLMs suffer from hallucinations affecting long-horizon task planning. In addition, [17] assumes correct execution of each step and does not account for occasional failures or unforseen changes of the environment. Our approach mitigates these issues using multiple specialized agents and intermediate environment feedback.
@@ II-C LLM-Based Multi-Agents in Robotics
Recently, numerous studies have focused on using LLMs for sequential decision-making [37, 38, 39] and complex reasoning tasks [40]. There is also an emerging field of using multiple agents driven by LLMs to solve complex tasks including robotics in a collaborative manner [41, 42, 43]. For most of the work in robotics, the LLM is used in multi-robot collaboration and communication. Additionally, all these work use RL policy agents which causes limits their generalization to new tasks and environments [44]. In contrast, we propose a multi-agent framework, MALMM, which uses three LLM agents in a zero-shot manner to perform robotics manipulation tasks.
@@ III METHOD
Large Language Models have recently emerged as a universal tool for a variety of tasks including robotic navigation and manipulation [45, 36, 34, 18, 17]. Such models exhibit surprising generalization abilities and support zero-shot solutions for new tasks. While several recent methods explored LLMs for high-level task planning [32, 31, 30, 36], other methods attempt to bring the power of LLMs all the way to the low-level trajectory planning [18, 17]. We follow this work and leverage LLMs for precise control of the end effector to solve complex and previously unseen manipulation tasks.
Trajectory planning is a complex problem that requires reasoning about the precise shape, position and semantics of the objects as well as understanding object interactions. To solve this problem, LLM agents can be instructed to handle information about the objects and the robot by text prompts. To harness the complexity of the problem, we here propose to address the manipulation problem with multiple LLM agents. Inspired by recent work on multi-agent LLMs [46], we design agents specialized on different aspects of the manipulation problem and connect them into a network. Below, we describe our proposed Multi-Agent Large Language Model for Manipulation (MALMM) framework (Fig. 2) in detail.
@@ III-A Preliminaries
Our goal is to design a robotics system capable of solving previously unseen manipulation tasks defined by natural language instructions. We assume access to a robotics environment supporting manipulation actions of closing the gripper, opening the gripper, and changing the gripper’s position and orientation. Observations of the environment are given either in the form of object and gripper 3D poses obtained from the simulator, or a 3D point cloud obtained from a depth camera. In the latter case, we estimate object poses and potential grasping poses from the point cloud. We also assume to have access to a modern LLM that can process language instructions together with environment observations in the form of text.
@@ III-B MALMM: Multi-Agent Manipulation
The core motivation of our work is to investigate how a multi-agent framework can leverage the collaborative and role-specific reasoning capabilities of multiple LLM agents to complete complex manipulation tasks. Below, we first introduce a single-agent architecture and then propose its multi-agent extensions. Single- and multi-agent architectures considered in our work are illustrated in Fig. 3.
Single Agent (SA). Prompting LLMs to interpret natural language instructions in the form of executable primitive actions has been shown successful for a variety of tasks including image understanding [47] and numerical reasoning [48]. In a similar fashion LLMs can be prompted to interpret embodied tasks, e.g., open a wine bottle and convert them into a sequence of primitive manipulation actions. A version of such a system with a single LLM agent is outlined in Fig. 3(a). Here a LLM is first prompted to break down the language instruction into a sequence of steps. It then uses its code generation capabilities [49] to translate each step into executable Python code, using predefined functions to control the end-effector. This code is then sent to a Python interpreter that executes the steps in the environment. After each step, LLM receives new observations from the environment and proceeds in a loop with planning and code generation until meeting termination criteria.
Multi-Agent (MA). Manipulation tasks require prolonged interactions with the environment. LLMs, however, often struggle to manage long context conversations [16], leading to hallucinations. To address this issue, we propose two specialized LLM agents with shorter role-specific contexts: the Planner and the Coder, see Fig. 3(b). The Planner breaks down the language instructions into a sequence of manipulation steps, while the Coder iteratively translates these steps into the executable Python code. After each intermediate step, the Planner detects potential failures and re-plans according to new observations of the environment.
Multi-Agent with Supervisor (MALMM). Our final multi-agent architecture (MALMM) extends MA with a Supervisor agent that coordinates the Planner, the Coder and the Code Executor, see Fig. 3(c). The Supervisor LLM decides which agent or tool to activate next based on input instructions, the roles of individual agents, environmental observations and the entire chat history of active agents.
@@ III-C Multi-Agent Prompting
Each agent is provided with a task-agnostic system prompt. The agents rely solely on their internal world knowledge for reasoning and decision-making. For prompt construction, we draw inspiration from [17] and its study of LLM based trajectory generation for robotics manipulation. Note, that in contrast to other recent work [18, 34] we do not provide agents with any examples for in-context learning and apply MALMM with no changes, i.e. in a zero-shot mode, to new tasks. Complete prompts used for LLM agents in our experiments are available from [50].
Each agent’s prompt is specifically designed to suit its role. Since LLMs require step-by-step reasoning to solve tasks, the Planner is prompted to generate steps that define the intermediate goals needed to complete the task. MALMM perceives the 3D environment using object poses provided either by the internal state of the simulator or by the analysis of visual observations, see Sec. III-D. Therefore, the Planner is given a detailed description of the environment’s coordinate system, enabling it to interpret directions from the gripper’s perspective. Given the limited exposure of LLMs to grounded physical interaction data in their training, LLM agents often fail to account for potential collisions. To address this, we include generic collision-avoidance rules in the Planner’s prompts. Additionally, to handle intermediate failures, primarily due to collisions or missed grasping, the Planner is prompted to evaluate outcomes after each intermediate step and, if necessary, replan based on updated observations. The prompt used for the Planner agent is shown in Figure 8.
@@
III-D Environmental Observations
LLMs trained on textual inputs cannot directly perceive or interpret 3D environments. Our agents receive information about the environment either from the internal state space of a simulator or from visual observations.
State space observations. In this setup LLM agents have direct access to the simulator state information. The observations are provided in the form of 3D bounding boxes, orientations and colors of scene objects as well as the 3D position, orientation and the open/close state of the gripper. The execution logs for empty container task presented in AppendixLABEL:app:output show input observation format as CURRENT ENVIRONMENT STATE for MALMM. The same format is also used for Single Agent and Multi-Agent presented in Section III-B.
Visual observations. To approach real-world settings, we restrict observations to images and 3D point clouds obtained from RGB-D sensor and then use pretrained vision models to extract information about scene objects. To this end, we deploy LangSAM [51] to generate image segmentation masks for objects described by text, e.g. ”block” or ”red jar” in task instruction. We then segment 3D object point clouds by projecting 2D segmentation masks into 3D. To compute accurate object-centric grasping poses, we apply M2T2 [52] model to predict grasp poses for 3D object point clouds for the object identified for grasping, see details in Appendix-F. We use obtained gripper poses to control the robotics gripper during grasping. To facilitate object placement, we estimate the 3D bounding box of the target object in the environment based on the task instructions. For example, in the task close the red jar, the target object would be the red jar. We extract the 3D bounding box directly from the object’s point cloud and use it to guide the placement process. Please refer to Appendix-F for details on obtaining visual observations.
@@ IV EXPERIMENTS
| CAP [18] | 0.00 | 0.00 | 0.00 | 8.00 | 0.00 | 0.00 | 76.00 | 0.00 | 0.00 | 9.33 |
| VoxPoser [34] | 20.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 36.00 | 64.00 | 32.00 | 16.89 |
| Single Agent (SA) | 52.00 | 40.00 | 48.00 | 24.00 | 44.00 | 80.00 | 92.00 | 48.00 | 20.00 | 49.78 |
| MALMM | 88.00 | 84.00 | 64.00 | 68.00 | 92.00 | 96.00 | 100.00 | 80.00 | 56.00 | 80.89 |
Our experiments are designed to assess whether MALMM can accurately solve a diverse set of manipulation tasks in a zero-shot manner.
@@
We use gpt-4-turbo111Experiments were conducted in June 2024 using the latest checkpoint. [54] to drive each of the three LLM agents in MALMM. For developing the multi-agent framework, we used AutoGen [55], which is an open-source programming library for building AI agents and facilitates collaboration among multiple agents to solve tasks. To perform true zero-shot evaluation, we refrain from using in-context learning examples within any of our agents. Our prompts were initially developed for the Stack Blocks task and are used for other tasks without any task specific tuning. MALMM generates 3D waypoints, while trajectories are computed and executed using a motion planner, following the approach commonly used in RLBench.
@@
The simulation is conducted in CoppeliaSim, interfaced through PyRep, using a Franka Panda robot with a parallel gripper. The setup incorporates RLBench [53], a robot learning benchmark that provides various language-conditioned tasks with specified success criteria. For our evaluation we sample 9 RLBench tasks, which feature diverse object shapes and task horizons. Fig. 4 shows snapshots of the 9 considered tasks.
In our state-space observation setup, the input observations are captured as bounding box coordinates and object orientations relative to the simulator’s ground truth reference axes. To obtain visual observations, we use single camera with the frontal view of the scene, see Figure 4. The point cloud is derived from the depth map corresponding to this view.
@@
We compare our approach against two state-of-the-art zero-shot LLM-based manipulation methods: Code as Policies (CAP) [18], and VoxPoser [34]. For CAP222https://code-as-policies.github.io/, we build upon its official PyBullet [56] implementation and adapt it for the use with RLBench. The official implementation of VoxPoser333https://voxposer.github.io/ was used without any modifications. Our Single-Agent baseline is a single LLM with feedback from the environment after the completion of each intermediate step. Additionally, for fair comparison between the Single Agent (SA) and MALMM, the prompts provided to both of them are equivalent. The only difference is that, in the case of MALMM, the information and the instructions in the Single-Agent prompt are distributed among the Supervisor, the Planner, and the Coder according to their respective roles. The prompts for Single Agent and MALMM are provided in Appendix-C1 and Appendix-C2 respectively.

### v2

# Multi-Agent Large Language Models for Zero-Shot Robotic Manipulation

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable planning abilities across various domains, including robotic manipulation and navigation. While recent work in robotics deploys LLMs for high-level and low-level planning, existing methods often face challenges with failure recovery and suffer from hallucinations in long-horizon tasks. To address these limitations, we propose a novel multi-agent LLM framework, Multi-Agent Large Language Model for Manipulation (MALMM). Notably, MALMM distributes planning across three specialized LLM agents, namely high-level planning agent, low-level control agent, and a supervisor agent. Moreover, by incorporating environment observations after each step, our framework effectively handles intermediate failures and enables adaptive re-planning. Unlike existing methods, MALMM does not rely on pre-trained skill policies or in-context learning examples and generalizes to unseen tasks. In our experiments, MALMM demonstrates excellent performance in solving previously unseen long-horizon manipulation tasks, and outperforms existing zero-shot LLM-based methods in RLBench by a large margin. Experiments with the Franka robot arm further validate our approach in real-world settings.
@@ I INTRODUCTION
Robotic manipulation has seen impressive advancements, enabling agents to handle increasingly complex tasks with greater precision and efficacy. Current solutions, however, often struggle with generalization, in particular when using imitation learning for policy training [1, 2]. Such methods typically excel at specific tasks but lack the adaptability to handle new tasks. One major drawback of imitation learning is the labor-intensive and time-consuming process for data collection, which limits the scalability of resulting policies. Moreover, training task-specific manipulation policies typically require thousands of training episodes [1, 2], making the approach computationally expensive and inefficient. To cope with the generalization, robotics policies should demonstrate a deeper understanding of their environment. This involves recognizing and grounding relevant objects and understanding the relationships between them [3]. Equipped with this knowledge, policies can then plan and execute actions more efficiently while adapting to changes in the environment and new task requirements.
@@
Recent advancements in LLMs have demonstrated remarkable generalization and reasoning capabilities across diverse domains such as commonsense, mathematical, and symbolic reasoning [4]. These models, particularly when scaled to billions of parameters, exhibit emergent abilities to break down complex tasks into simpler steps through structured reasoning techniques such as chain of thought [5]. LLMs have already shown promise in high-level task planning in different domains, suggesting their potential for flexible and versatile robotic manipulation. Yet, recent LLM-based methods for robotic planning face multiple challenges. One important issue is the tendency of LLMs to produce incorrect high-level plans and low-level control. Additionally, LLMs suffer from hallucinations in long-context generation [6], which is often observed in closed-loop LLM systems. As a result, they may disregard geometric constraints and the parameters of predefined functions or may even lose the sight of the goal.
@@
In this work, we propose a Multi-Agent Large Language Model for Manipulation (MALMM) to leverage the collective intelligence and the specialized skills of multiple agents for complex manipulation tasks. Our framework incorporates agents dedicated to high-level planning, low-level code generation and a supervisor that oversees transitions between other agents and tools. We show that through the use of multiple specialized agents in a multi-agent setup, we are able to mitigate the hallucination issues observed in the case of a single agent, as shown in Fig. 1.
Our contributions can be summarized as follows:
We introduce the first multi-agent LLM framework for robotic manipulation MALMM, equipped with specialized agents that bring collaborative and role-specific capabilities to handle unseen, diverse, and complex manipulation tasks.
We demonstrate the advantages of the proposed multi-agent framework through systematic ablation studies on tasks with varying horizons and complexity.
We evaluate MALMM in challenging zero-shot settings both in a simulation and in the real world, and we show substantial improvements over state-of-the-art methods.
@@ II-A Language Grounded Robotics
Language instructions enable the definition of complex robotics tasks with compositional goals [7] and support scalable generalization to new tasks [8]. The literature around language grounding is vast, ranging from classical tools such as lexical analysis, formal logic, and graphical models to interpreting language instructions [7, 9]. Recently, much effort has focused on adopting the impressive capabilities of LLMs to language grounding in robotics [8]. Additionally, recent advancements have benefited from pre-trained LLMs thanks to their open-world knowledge, tackling more challenging tasks such as 3D robotic manipulation and leveraging code generation capabilities to produce high-level, semantically-rich procedures for robot control.
@@ II-B LLM for Robotics
Language Models have been used for various robotics purposes including the definition of reward functions [10], task planning [11, 12, 13], failure summarization and guiding language-based planners [14], and policy program synthesis [15]. VoxPoser [16] and Language to Rewards [17] used LLMs for generating reward regions for assisting external trajectory optimizers in computing trajectory. Our work is most related to methods using LLMs for manipulation planning. Most of such work [12, 16] relies on pre-trained skills, motion primitives, and trajectory optimizers and has focused primarily on high-level planning. The closest to our approach is Language Models as Zero-Shot Trajectory Generators [11], which deployed a language model to generate high-level plans and then convert these plans into low-level control. However, LLMs suffer from hallucinations, which affect long-horizon task planning. Moreover, [11] assumed the correct execution of each step and did not account for occasional failures or unforeseen changes in the environment. Our evaluation shows sizable improvements of MALMM over [11] thanks to its multiple specialized agents and intermediate environment feedback.
@@ II-C LLM-Based Multi-Agents in Robotics
Recently, several studies have focused on using LLMs for sequential decision-making and complex reasoning tasks [18]. There is also an emerging field of using multiple agents driven by LLMs to solve complex tasks, including robotic manipulation, in a collaborative manner [19]. In most of the work in robotics, LLMs have been used in multi-robot collaboration and communication. Moreover, Reinforcement Learning (RL) policy agents are also used in collaboration with LLMs, which limits their generalization to new tasks and environments [20]. In contrast, we propose a multi-agent framework, MALMM, which incorporates three role-specific LLM agents and enables zero-shot execution of previously unseen robotic manipulation tasks.
@@ III METHOD
LLMs have recently emerged as a universal tool for a variety of tasks, including robotic navigation and manipulation [21, 16, 15, 11]. Such models exhibit surprising generalization abilities and support zero-shot solutions for new tasks. While several recent methods have explored LLMs for high-level task planning [22, 12], other methods attempted to bring the power of LLMs to the low-level trajectory planning [15, 11]. We follow these works and leverage LLMs to precisely control the end-effector for solving complex and previously unseen manipulation tasks, without relying on any predefined skill library for sub-tasks in the manipulation process.
Trajectory planning is a complex problem that requires reasoning about the precise shape, position, and semantics of the objects as well as understanding object interactions. To solve this problem, LLM agents can be instructed to handle information about objects and robots by text prompts. To harness the complexity of the problem, here we propose to address the manipulation problem with multiple LLM agents. Inspired by recent work on multi-agent LLMs, we design agents specialized in different aspects of the manipulation problem and connect them into a framework. Below, we describe our proposed MALMM framework in detail (see Fig. 2).
@@ III-A Preliminaries
Our goal is to design a LLM framework capable of solving previously unseen manipulation tasks defined by natural language instructions. We assume access to a robotic environment supporting manipulation actions of closing the gripper, opening the gripper, and changing the gripper’s position and orientation around the z axis. Environment observations are obtained as described in Sec. III-D.
@@ III-B MALMM: Multi-Agent Manipulation
The core motivation of our work is to investigate how a multi-agent framework can leverage the collaborative and role-specific reasoning capabilities of multiple LLM agents to complete complex manipulation tasks. Below, we first introduce a single-agent architecture and then we propose its multi-agent extensions. The single and multi-agent architectures considered in our work are illustrated in Fig. 3.
Single Agent (SA). SA is adopted from [11] by environment-specific (RLBench [23] or real-world) prompt tuning. Prompting LLMs to interpret natural language instructions in the form of executable primitive actions has been shown successful for a variety of tasks, including image understanding [24] and numerical reasoning [25]. Similarly, LLMs can be prompted to interpret embodied tasks, e.g., open a wine bottle and convert them into a sequence of primitive manipulation actions. A version of such a system with a single LLM agent is outlined in Fig. 3(a). Here, an LLM is first prompted to break down the language instruction into a sequence of steps. It then uses its code generation capabilities to translate each step into executable Python code, using predefined functions to control the end-effector. This code is then sent to a Python interpreter that executes the steps in the environment. After each step, the LLM receives new observations from the environment and proceeds in a loop with planning and code generation until meeting termination criteria.
Multi-Agent (MA). A Single Agent performing multiple roles struggles to excel in all of them. To address this issue, we propose two specialized LLM agents with shorter role-specific contexts: the Planner and the Coder, see Fig. 3(b). The Planner breaks down the language instructions into a sequence of manipulation steps while the Coder iteratively translates these steps into an executable Python code. After each intermediate step, the Planner detects potential failures and re-plans according to new observations of the environment.
Multi-Agent with a Supervisor. Our final multi-agent architecture, MALMM extends MA with a Supervisor agent that coordinates the Planner, the Coder, and the Code Executor, as shown in Fig. 3(c). The Supervisor decides which agent or tool to activate next based on the input instructions, the roles of the individual agents, the environment observations, and the entire chat history of the active agents.
@@ III-C Multi-Agent Prompting
Each agent is provided with a task-agnostic system prompt. The agents rely solely on their internal world knowledge for reasoning and decision-making. For prompt construction, we draw inspiration from [11] and its study of LLM-based trajectory generation for robotic manipulation. We adapt the prompt according to our environments (RLBench and real-world). Note that unlike other recent work [15, 16], we do not provide the agents with any examples for in-context learning and apply MALMM to new tasks without any changes, i.e., in a zero-shot mode.
Each agent’s prompt is specifically designed to suit its role. Since LLMs require step-by-step reasoning to solve tasks, the Planner is prompted to generate steps that define the intermediate goals needed to complete the task. MALMM perceives the environment observations directly from the simulator or by analyzing the RGBD (see Sec. III-D). Therefore, the Planner is given a detailed description of the environment’s coordinate system, enabling it to interpret the directions from the gripper’s perspective. Given the limited exposure of LLMs to grounded physical interaction data in their training, LLM agents often fail to account for potential collisions. To address this, we include generic collision-avoidance rules in the Planner prompts. Moreover, to handle intermediate failures, primarily due to collisions or missed grasps, we prompt the Planner to evaluate action based on the previous and current environment observations after each intermediate step, and to replan if necessary. The prompt used for the Planner agent is shown in Figure 10.
@@
III-D Environment Observations
LLMs trained on textual inputs cannot directly perceive or interpret 3D environments. Our agents receive information about the environment either from the internal state space of a simulator or from RGBD.
State-space observations. In this setup, the LLM agents have direct access to the simulator’s state information. The observations are provided as 3D bounding boxes (object dimensions, center position, and orientation), along with object colors and the gripper’s position, orientation, and open/closed state. The execution logs for empty container task presented in AppendixLABEL:app:output show input observation format as CURRENT ENVIRONMENT STATE for MALMM. The same format is also used for Single Agent and Multi-Agent presented in Section III-B.
Visual observations. To apply MALMM in real-world settings, we restrict observations to the front-facing RGB-D sensor (as illustrated in Fig. 4), from which we obtain both RGB images and 3D point clouds. We then use pre-trained foundation models to extract information about scene objects. To this end, we employ gpt-4-turbo to derive a list of objects relevant to the instruction text and RGB image, and subsequently utilize LangSAM [26] to produce segmentation masks for these objects, such as block or red jar. We then segment the 3D object point clouds by projecting the 2D segmentation masks into the 3D space. To compute accurate object-centric grasping poses, we apply the M2T2 [27] model and predict grasps given 3D point clouds of source objects. We use the obtained gripper poses to control the gripper during grasping. To facilitate object placement, we estimate the 3D bounding box of the target object in the environment. For example, in the task close the red jar, the target object would be the red jar. We extract the 3D bounding box directly from the object’s point cloud, and use it to guide the placement process. We leverage these visual observations while conducting experiments in both simulated and real-world settings. Please refer to Appendix-F for details on obtaining visual observations.
@@ IV EXPERIMENTS
| CAP [15] | 0.00 | 0.00 | 0.00 | 0.08 | 0.00 | 0.00 | 0.76 | 0.00 | 0.00 | 0.09 |
| VoxPoser [16] | 0.20 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.36 | 0.64 | 0.32 | 0.17 |
| Single Agent (SA) [11] | 0.52 | 0.40 | 0.36 | 0.24 | 0.44 | 0.80 | 0.92 | 0.48 | 0.20 | 0.50 |
| MALMM (†) | 0.84 | 0.88 | 0.60 | 0.80 | 0.64 | 0.84 | 0.84 | 0.56 | 0.32 | 0.70 |
| MALMM | 0.88 | 0.84 | 0.64 | 0.68 | 0.92 | 0.96 | 1.00 | 0.80 | 0.56 | 0.81 |
We evaluate the accuracy of MALMM in a zero-shot settings, i.e. when solving diverse set of previously unseen tasks defined by a short text description.
@@
We use gpt-4-turbo111The experiments were conducted in June 2024 using the gpt-4-turbo-2024-04-09 snapshot. [28] to drive the LLM agents in all our experiments. Additionally, we report the results of MALMM using LLaMA-3.3-70B [29] to demonstrate the performance of our framework with an open-source LLM model. For developing the multi-agent framework, we used AutoGen [30], which is an open-source programming library for building AI agents and facilitates collaboration between multiple agents to solve complex tasks. To perform zero-shot evaluation, we do not fine-tune our agents, and we use no training data for in-context learning. We initially developed our prompts for the Stack Blocks task and used them for other tasks without any task-specific tuning. MALMM generates 3D waypoints, while the trajectories are computed and executed using a motion planner, following the approach commonly used in RLBench. Our code, prompts, and additional results are available from the project webpage [31].
@@
We conduct the simulation in CoppeliaSim, interfaced via PyRep, using Franka Panda robot with a parallel gripper. The setup incorporates RLBench [23], a robot learning benchmark that provides various language-conditioned tasks with specified success criteria. For evaluation, we sampled 9 RLBench tasks with 25 samples per task, featuring diverse object poses, object shapes, and task horizons. Fig. 4 shows snapshots of the nine considered tasks.
In the real-world setup, we evaluated five tasks on a tabletop using a 7-DOF Franka Emika Panda Research 3 robot equipped with a parallel jaw gripper. Three of these tasks are identical to those that use vision-based observations from the simulator, while two are new. We use an Intel RealSense D435i RGB-D camera to capture the frontal view and the panda-py [32] library to control the robot arm.
@@
We compare our approach to three state-of-the-art zero-shot LLM-based manipulation methods: Code as Policies (CAP) [15], VoxPoser [16], and SA [11]. For CAP222https://code-as-policies.github.io/, we adapt its official implementation for use with RLBench. We used the official implementation of VoxPoser333https://voxposer.github.io/ without any modifications. Our SA baseline is a version of “Language Models as Zero-Shot Trajectory Generators” [11] with environment-specific (RLBench or real-world) prompt tuning. Additionally, for fair comparison between the SA and MALMM, the prompts we provided to both of them were equivalent. The only difference is that in the case of MALMM, the information and the instructions in the Single-Agent prompt are distributed between the Supervisor, the Planner, and the Coder according to their respective roles. All the baselines use gpt-4-turbo as an LLM. The prompts for Single Agent and MALMM are provided in Appendix-C1 and Appendix-C2 respectively.